OpenStudy (anonymous):

a necklace hanging from two fixed points A and B at the same level. the length of the necklaces between two points is 100cm. the midpoint of the necklace is 8 cm below A and B. assume that the necklaces hang in parabolic curve , find the equation of the curve ( solution )

hartnn (hartnn):

Let us assume that the necklace is hanging like shown below in the figure, the two points where the parabola(or necklace)meets the x axis is kept (0,0) and(100,0)to ensure distance as 100 between A and B.Then two points certainly lie on that parabola,as shown which are (50,-8) and (100,0).....this we put in standard parabola equation:y=a*x^2+x*b Two coordinates; x = 50, y = -8 50^2a + 50b = -8 2500a + 50b = -8 and x = 100, y = 0 100^2a + 100b = 0 10000a + 100b = 0 : Multiply the 1st equation by 2, subtract from the 2nd equation 10000a + 100b = 0 5000a + 100b = -16 ---------------------subtraction eliminates b find a 5000a = 16 a = 16/5000 a = 0.0032 : Find b 2500(0.0032) + 50b = -8 8 + 50b = -8 50b = -8 - 8 b = -16/50 b = -0.32 which finally gives the equation of parabola as y = 0.0032x^2 - 0.32x

OpenStudy (anonymous):

thanks...

hartnn (hartnn):

welcome,but did u understand the solution fully?ask if there are still doubts...